find the exact roots for the following: x^3-8x^2+4x+48
Answer by ewatrrr(24785)


Hi
finding the exact roots
x^3-8x^2+4x+48 = 0 |substitution show x = -2 is a solution
(x+2)(x^2-10x + 24) |Dividing x^3-8x^2+4x+48 by (x+2)
factoring quadratic
(x+2)(x-6)(x-4) = 0
Note:SUM of the inner product(-6x) and the outer product(-4x) = -10x
(x+2)(x-6)(x-4) = 0
x = {-2,6,4}